# Effective and Transparent RAG: Adaptive-Reward Reinforcement Learning for Decision Traceability

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has significantly improved the performance of large language models (LLMs) on knowledge-intensive domains. However, although RAG achieved successes across distinct domains, there are still some unsolved challenges: 1) Effectiveness. Existing research mainly focuses on developing more powerful RAG retrievers, but how to enhance the generator’s (LLM’s) ability to utilize the retrieved information for reasoning and generation? 2) Transparency. Most RAG methods ignore which retrieved content actually contributes to the reasoning process, resulting in a lack of interpretability and visibility. To address this, we propose ARENA (Adaptive-Rewarded Evidence Navigation Agent), a transparent RAG generator framework trained via reinforcement learning (RL) with our proposed rewards. Based on the structured generation and adaptive reward calculation, our RL-based training enables the model to identify key evidence, perform structured reasoning, and generate answers with interpretable decision traces. Applied to Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, abundant experiments with various RAG baselines demonstrate that our model achieves 10–30% improvements on all multi-hop QA datasets, which is comparable with the SOTA Commercially-developed LLMs (e.g., OpenAI-o1, DeepSeek-R1). Further analyses show that ARENA has strong flexibility to be adopted on new datasets without extra training. Our models and codes are publicly released [^1] https://github.com/ren258/ARENA.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful paradigm for enhancing large language models (LLMs) with non-parametric knowledge [8, 12, 25, 55], particularly in knowledge-intensive domains such as medicine [23], law [49], and finance [37]. While retrieved information can significantly improve generation performance [22, 52], existing research has mainly concentrated on improving retrievers, with relatively less attention given to the generator component. Our empirical analysis reveals a notable performance gap: on widely-used and challenging multi-hop QA benchmarks, 7B-scale models using identical retrieved contexts underperform state-of-the-art reasoning models like OpenAI-o1 [30] and DeepSeek-R1 [42] by 15–35% in answer accuracy (see Appendix A). This indicates that the generator’s reasoning capability, not just the quality of retrieval, remains a key bottleneck in current RAG pipelines.
Beyond boosting effectiveness, advancing the generator opens up valuable opportunities for improving transparency — critical qualities for deploying LLMs in real-world, high-stakes applications to understand which retrieved document is involved in reasoning and generation [41, 48, 24]. Many generators produce unstructured answers without exposing the underlying decision process, limiting their interpretability and trustworthiness. The central challenge lies in enabling the generator to identify, integrate, and reason over relevant evidence from noisy or redundant contexts [13]. Smaller open-source models face greater difficulty in such scenarios due to limited reasoning capacity [9].
Recently, reinforcement learning (RL) has shown promise in improving the reasoning ability of LLMs, especially in math [53] and code [26] generation, which often rely on reward signals derived from structured task objectives, enabling models to learn reasoning behavior beyond what supervised fine-tuning can offer. This inspires us to explore RL as a path for improving RAG generation. Existing RL-based approaches—such as GRPO [11]—typically use generic reward designs and output formats, which fail to capture the structure of multi-hop QA [34]. Moreover, unstable KL regularization often leads to divergence during training, limiting their applicability in retrieval-grounded tasks.
To address these challenges, we propose ARENA (Adaptive-Rewarded Evidence Navigation Agent), a transparent and effective RAG generator framework trained via reinforcement learning. ARENA introduces a structured output format that includes selected references, explicit reasoning traces, and final answers, enabling end-to-end interpretability. We further design a suite of adaptive, task-specific rewards and a stabilized optimization strategy tailored for multi-hop QA. Applied to open-source models like Qwen2.5-7B-Instruct [50] and Llama3.1-8B-Instruct [10], ARENA achieves 10–30% accuracy improvements across three datasets, rivaling the best closed-source systems while offering better transparency and generalizability. Our main contributions are listed as follows:
  • We identify two key challenges in current RAG systems, namely limited reasoning ability and lack of interpretability, and call for greater attention to the generator optimization.
  • We propose ARENA, a reinforcement learning-based framework for RAG that integrates structured generation, adaptive reward design, and stabilized training for transparent and effective RAG reasoning.
  • Experiments demonstrate that ARENA outperforms all open-source baselines of comparable scale, achieving results on par with leading closed-source models across three challenging multi-hop QA datasets. ARENA also generalizes well across datasets and backbones.
  • All the codes and models are released to support future study on optimizing RAG generator.
@@ 2 Related Work
2.1 Retrieval-Augmented Generation
RAG has emerged as an effective approach to enhancing model performance across a wide range of NLP tasks by incorporating externally retrieved documents into the input [14, 8], particularly for knowledge-intensive applications such as open-domain question answering (ODQA) [22]. By grounding the model’s generation in retrieved evidence, RAG can substantially mitigate hallucinations and factual inaccuracies [40, 2]. Moreover, recent studies have demonstrated that the benefits of retrieval become increasingly pronounced as model scale grows [28, 39], further underscoring the practical value of retrieval-based methods. In the ODQA task, the prevailing paradigm follows a retriever-and-reader framework [1], where significant effort has been devoted to enhancing the retriever. Approaches include decomposing complex queries into sub-questions using prompts [33], rewriting queries to enhance retrieval quality [27], guiding retrieval through CoT reasoning [46], constructing summaries to select the most plausible answer among candidates [20], and compressing prompts while integrating retrieval information [18]. Despite these substantial advances, how to optimize the RAG generator was ignored by most studies.
The multi-hop task [51, 15, 45, 33] poses significant challenges in the field of RAG, as they require LLMs to integrate information across multiple retrieved documents in order to generate a correct answer, which requires more powerful reasoning and generation ability. However, the presence of irrelevant or noisy contexts can adversely affect answer generation, resulting in factual errors [32, 6]. Thus, strengthening the reader’s capability to reason over and selectively utilize retrieved content is critical for achieving robust performance.
Beside, some studies have demonstrated that verifying the reliability of LLM-generated content based on citations is an effective approach to mitigating hallucinations and provides transparent RAG for users [17]. But most of these studies are post-hoc based, which means the citations are not generated during LLM reasoning. We propose that, this step can also be finished by the RAG generator.
2.2 LLMs with RL
RL serves as a powerful tool for LLMs to improve their alignment to human values and reasoning abilities. In particular, reinforcement learning with human feedback (RLHF) has demonstrated the effectiveness of aligning model outputs with human preferences [31]. Beyond alignment, models such as OpenAI-o1 [30], DeepSeek-R1 [11] and Kimi k1.5 [44] showcase significant gains in mathematical reasoning and code generation following RL-based training. More recently, R1-Searcher [42] and ReSearch [4] have leveraged Reinforce++ [16] and GRPO [38], respectively, to enhance the abilities of the models to interact with retrieval systems and produce accurate responses. Collectively, these methods highlight how reinforcement learning reward signals can systematically improve the performance of LLM in unsupervised settings. However, prior methods have primarily focused on strengthening general reasoning ability, without explicitly designing reward functions tailored to specific tasks such as RAG. In this work, we introduce specialized incentives to address this gap.
@@ 2.2 LLMs with RL
Figure 1: Overview of the ARENA (Adaptive-Rewarded Evidence Navigation Agent) framework. The system is composed of three key components: ① Structured Generation, where the policy model generates a multi-part response including selected evidence, reasoning analysis, and a final answer; ② KL Stabilization, we modify the KL formulation to improve stability; ③ Adaptive Reward Calculation, where the model outputs are evaluated on four axes—format, accuracy, relevance, and bonus—to provide interpretable and fine-grained training signals.
@@ 3 ARENA: Adaptive-Rewarded Evidence Navigation Agent
We introduce ARENA (Adaptive-Rewarded Evidence Navigation Agent), a novel reinforcement learning framework for the training of RAG generators. ARENA enhances the reasoning capabilities of language models and improves decision transparency, enabling more interpretable and reliable deployment in knowledge-intensive domains. As shown in Figure 1, ARENA consists of three key components: (1) Structured generation, where the model produces evidence-grounded multi-part responses; (2) KL stabilization, we stabilized KL regularization for policy refinement; and (3) Adaptive reward calculation, which provides interpretable feedback across multiple dimensions. We will introduce these modules in detail in the following subsections.
@@ 3.1 Framework formalization
Table 1: ARENA prompt template format. The <relevance> part is the key to encourage explicit and auditable evidence selection. The placeholders {question} and {references} are filled in at runtime.
@@ 3.2 Reinforcement learning with stability enhancements
In complex reasoning tasks, supervised fine-tuning (SFT) often falls short due to the lack of high-quality, annotated reasoning traces [5]. Recently, reinforcement learning (RL) has emerged as a powerful alternative to enhanceperformance through self-improvement. Motivated by this, ARENA employs reinforcement learning to activate and refine the reasoning capabilities of instruction-tuned models to enhance the RAG generation ablities of them.
@@
GRPO [38] is an advanced policy optimization algorithm that improves over Proximal Policy Optimization (PPO) [36] by eliminating the need for a critic model. For each training instance $q$, the policy $\pi_{\theta}$ generates $G$ outputs $\{o_{i}\}_{i=1}^{G}$. Each output is scored using the task-specific reward function (detailed in Section 3.3), and a group-wise normalized advantage is computed as: $\mathcal{A}_{i}=\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},\text{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\epsilon,1+\epsilon\right)\right)\frac{r_{i}-{\mathrm{m}ean(\{r_{1},r_{2},\cdots,r_{G}\})}}{{\mathrm{s}td(\{r_{1},r_{2},\cdots,r_{G}\})}}$, where $r_{i}=\sum_{j=1}^{N}w_{j}\cdot R_{j}(o_{i}\mid q)$ is the weighted reward for output $o_{i}$.
The KL divergence between the current policy and a fixed reference model is incorporated as a regularization term to prevent policy collapse. The final GRPO objective is:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\mathcal{A}_{i}-\beta\,\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{ref}\right)\right)\right].$
@@
The KL divergence term plays a key role in constraining the updated policy from diverging excessively from the reference model. In standard GRPO, it is estimated using the following unbiased formulation:
$r=\frac{\pi_{\mathrm{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)},\quad\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{\mathrm{ref}})=r-\log r-1,$
which is theoretically appealing: it is non-negative, unbiased, and has low variance under common assumptions. However, this estimator becomes numerically unstable when the policy assigns near-zero probability to sampled outputs ($\pi_{\theta}(o\mid q)\to 0$), causing the importance ratio $r\to\infty$. This results in gradient explosions and erratic updates that severely degrade training stability.
To address this, we aim to retain the desirable properties of the original KL estimator—namely, non-negativity, low variance, and smooth penalization—while ensuring numerical stability in rare-event sampling. Motivated by this, we use a simple but effective replacement [35]:
$\text{KL}_{\mathrm{stable}}(o)=\frac{1}{2}(\log r)^{2},$
which maintains non-negativity and provides a symmetric, smooth penalty for large divergences. It avoids singularities even when $\pi_{\theta}(o\mid q)$ is small, thus offering more stable gradients in training.
@@ 3.3 Adaptive reward design
Outputs that match the expected structural format—consisting of <relevance>, <analysis>, and <answer> parts in order—receive a reward of 1; others receive 0.

### v2

# Transparent and Robust RAG: Adaptive-Reward Reinforcement Learning for Decision Traceability

@@ Abstract
Retrieval-Augmented Generation (RAG) delivers substantial value in knowledge-intensive applications. Many recent works use reinforcement learning (RL) to effectively elicit strong reasoning in RAG generators. However, two key challenges remain unresolved: (1) Transparency: most prior methods do not explicitly indicate which references are actually used during the reasoning that leads to the final answer, limiting interpretability and visibility; (2) Stability: the KL divergence estimator used in existing RL-based approaches may cause gradient spikes, which can lead to unstable training. To address these challenges, we propose Adaptive-Rewarded Evidence Navigation Agent (ARENA), a transparent and robust RAG generator framework trained via RL with designed rewards. Based on our proposed structured protocol, KL divergence stabilization, and adaptive reward calculation modules, ARENA enables RAG generator to identify key evidence, perform structured reasoning, and generate answers with interpretable decision traces. Applied to Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, abundant experiments with various baselines demonstrate that our model achieves highly transparent outputs with 10–30% accuracy improvements across three multi-hop QA datasets, which is comparable with advanced closed-source LLMs (e.g., OpenAI-o1, DeepSeek-R1). Further analyses show ARENA has strong generalization to unseen datasets and tasks. Our models and codes are publicly released [^1] https://github.com/ren258/ARENA.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful paradigm for enhancing large language models (LLMs) with non-parametric knowledge (Gao et al., 2023; Gupta et al., 2024; Li et al., 2025b; Zhou et al., 2025), particularly in knowledge-intensive domains such as medicine (Li et al., 2024a), law (Wiratunga et al., 2024), and finance (Setty et al., 2024). Beyond accuracy gains from retrieved context (Lewis et al., 2020; Yu, 2022), practical deployment further requires transparency, auditability, and decision traceability, so that one can verify which evidence supports which reasoning steps and the final answer (Singh et al., 2024; Wei Jie et al., 2024; Li et al., 2024b).
Despite rapid progress, important gaps remain. Recent RL-based work emphasizes active acquisition of information, interleaving search and reasoning when initial documents are insufficient (Gao et al., 2025; Li et al., 2025a). However, two major challenges are still unresolved. Transparency: open-source systems generally do not make explicit or verifiable which retrieved references are actually used during the reasoning that leads to the final answer (Song et al., 2025a; Chen et al., 2025b), while strong closed-source commercial models provide citation without disclosing training or data acquisition strategies (Team, 2025; OpenAI, 2025); moreover, the literature lacks a quantitative assessment of citation relevance accuracy (Jia et al., 2024; Yu et al., 2023; Wei et al., 2024). Stability: current RL-based methods for RAG largely adopt GRPO (Guo et al., 2025), which shown to elicit powerful reasoning in math (Zeng et al., 2025) and code (Liu et al., 2024), but the KL divergence estimator used in these methods may induce sharp gradient spikes during training, and there is little theoretical or empirical analysis of how the choice of estimator affects training stability and generation quality in RAG (Tang & Munos, 2025; Yu et al., 2025).
To address these challenges, we propose Adaptive-Rewarded Evidence Navigation Agent (ARENA), a transparent and robust RAG generator trained with reinforcement learning. First, ARENA enforces a structured output protocol that explicitly includes the selected references, step-wise reasoning traces, and the final answer, rendering the entire evidence–reasoning–answer chain auditable. Second, it selects a more numerically stable KL divergence estimator to mitigate gradient spikes during optimization. Third, it employs adaptive, task-specific reward computation aligned with multi-hop QA (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022), encouraging correct answers, faithful citation usage, and well-formed decision traces.Applied to open-source backbones such as Qwen2.5-7B-Instruct (Yang et al., 2024) and Llama3.1-8B-Instruct (Grattafiori et al., 2024), ARENA produces highly transparent outputs and delivers 10–30% accuracy improvements across three datasets while approaching advanced closed-source LLMs. Moreover, our generalization experiments demonstrate that ARENA transfers well across different tasks, backbones, and retrieval settings, our ablation studies further verify the effectiveness of each individual component.
Our main contributions are listed as follows:
  • We propose ARENA, a reinforcement learning framework for RAG generator that improves both transparency and robustness. It integrates structured output protocol for explicit evidence–reasoning–answer traceability, KL-divergence stabilization that mitigates gradient spikes, and adaptive task-specific rewards tailored for multi-hop QA.
  • We highlight the choice of KL divergence estimator will impact the stability and performance of RL in RAG, with both theoretical and empirical analysis.
  • We validate our approach on three multi-hop QA benchmarks (HotpotQA, 2Wiki, and MuSiQue), where ARENA delivers 10–30% improvements in EM, F1, and LLM-as-a-judge metrics while maintaining superior transparency, approaching the performance of closed-source commercial models and demonstrating the effectiveness of our framework.
@@ 2 Related Work
2.1 RAG Foundations and the Transparency Gap
RAG grounds model outputs in external evidence (Guu et al., 2020; Gao et al., 2023; Lewis et al., 2020), improving factual consistency on knowledge-intensive tasks such as open-domain QA (Shuster et al., 2021; Chen et al., 2024). Complex multi-hop QA tasks (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2022; Press et al., 2023) are particularly challenging for RAG, where irrelevant or noisy context can mislead generation and cause factual errors (Petroni et al., 2020; Creswell et al., 2023). To mitigate such risks and improve interpretability, prior work explores citation-based auditing and related techniques (Huang & Chang, 2024; Jia et al., 2024; Yu et al., 2023; Wei et al., 2024). However, despite these efforts, existing studies still lack explicit and readily evaluable citation labels that make clear which retrieved references support intermediate reasoning and the final answer.
2.2 Search-Centric RAG: Progress and Remaining Stability Issues
Strengthening retrieval has been shown to improve RAG performance, and its benefits tend to grow with model scale (Mallen et al., 2023; Shi et al., 2024). Early approaches rely on prompt-based techniques to enhance retrieval (Press et al., 2023; Ma et al., 2023; Trivedi et al., 2023; Kim et al., 2024). More recently, research has focused on settings where initially retrieved documents are insufficient to answer the query, prompting models to actively acquire additional information until they can infer the answer. Supervised finetuning approaches synthesize retrieval trajectories and apply multi-criteria filtering to construct high-quality training data (Wang et al., 2025b; Sun et al., 2025b), while reinforcement learning optimizes search policies with reward feedback through online real search (Song et al., 2025b; Chen et al., 2025b) or offline simulated search (Mei et al., 2025; Sun et al., 2025a). Methodologically, most RL-based methods adopt GRPO or related policy-gradient variants (Shao et al., 2024; Hu, 2025); prior studies further suggest that commonly used KL divergence estimators can make training unstable (Tang & Munos, 2025; Yu et al., 2025). This observation motivates us to explore more stable KL divergence estimators for RAG training.
@@ 2.2 Search-Centric RAG: Progress and Remaining Stability Issues
Figure 1: Overview of the Adaptive-Rewarded Evidence Navigation Agent (ARENA) framework. The system is composed of three key components: ① Structured Protocol, where the policy model generates a multi-part response including selected evidence, reasoning analysis, and a final answer; ② KL Stabilization, we modify the KL divergence estimator to improve stability; ③ Adaptive Reward Calculation, where the model outputs are evaluated on four axes—format, accuracy, relevance, and bonus—to provide interpretable and fine-grained training signals.
@@ 3 ARENA: Adaptive-Rewarded Evidence Navigation Agent
We introduce Adaptive-Rewarded Evidence Navigation Agent (ARENA), a novel reinforcement learning framework for the training of RAG generators. ARENA enhances the reasoning capabilities of LLMs and improves decision transparency, enabling more interpretable and reliable deployment in knowledge-intensive domains. As shown in Figure 1, ARENA consists of three key components: (1) Structured protocol, where the model produces evidence-grounded multi-part responses; (2) KL stabilization, we stabilized KL regularization for policy refinement; and (3) Adaptive reward calculation, which provides interpretable feedback across multiple dimensions. We will introduce these modules in detail in the following subsections.
@@ 3.1 Framework formalization
Table 1: ARENA prompt template. The <relevance> part is the key to encourage explicit and auditable evidence selection. The placeholders {question} and {references} are filled in at runtime.
@@ 3.2 Reinforcement learning with stability enhancements
In complex reasoning tasks, supervised fine-tuning (SFT) often falls short due to the lack of high-quality, annotated reasoning traces (Chu et al., 2025). Recently, reinforcement learning (RL) has emerged as a powerful alternative to enhance performance through self-improvement. Motivated by this, ARENA employs reinforcement learning to activate and refine the reasoning capabilities of instruction-tuned models to enhance the RAG generation ablities of them.
@@
GRPO (Shao et al., 2024) is an advanced policy optimization algorithm that improves over Proximal Policy Optimization (PPO) (Schulman et al., 2017) by eliminating the need for a critic model. For each training instance $q$, the policy $\pi_{\theta}$ generates $G$ outputs $\{o_{i}\}_{i=1}^{G}$. Each output is scored using the task-specific reward function (detailed in Section 3.3), and a group-wise normalized advantage is computed as: $\hat{A}_{i,t}=\frac{r_{i}-\mathrm{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\mathrm{std}(\{r_{1},r_{2},\cdots,r_{G}\})},\quad r_{i}=\sum_{j=1}^{N}w_{j}\cdot R_{j}(o_{i}\mid q),$ .
let $\ell_{i,t}(\theta)=\min\!\left[\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}\mid q,o_{i,<t})}\hat{A}_{i,t},\,\text{clip}\!\left(\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}\mid q,o_{i,<t})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]$,
The final GRPO objective is:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\big(\ell_{i,t}(\theta)-\beta\,\mathbb{D}_{KL}(\pi_{\theta}\,\|\,\pi_{ref})\big)\right].$
@@
In standard GRPO, the KL divergence is instantiated via the $k_{3}$ estimator (Schulman, 2020). Let
$r(q,o_{i})\;=\;\frac{\pi_{\mathrm{ref}}(o_{i,<t}\mid q)}{\pi_{\theta}(o_{i,<t}\mid q)}\,.$
The $k_{3}$ estimate of the KL is
$L_{k_{3}}(\theta)\;=\;\mathbb{E}_{q,o_{i}}\!\left[\,r(q,o_{i})\;-\;\log r(q,o_{i})\;-\;1\,\right].$
In this work we instead adopt the $k_{2}$ estimator:
$L_{k_{2}}(\theta)\;=\;\frac{1}{2}\,\mathbb{E}_{q,o_{i}}\!\left[\big(\log r(q,o_{i})\big)^{2}\right].$
Compared with $k_{3}$, the $k_{2}$ estimator preserves non-negativity and, crucially, offers lower variance and gradient equivalence in expectation to the KL gradient; moreover, it provides symmetric penalties around r=1r{=}1 and exhibits improved stability when $\pi_{\theta}$ deviates from $\pi_{\mathrm{ref}}$. For these reasons, we use $k_{2}$ in our GRPO training. Detailed theoretical analysis and proof can be found in Appendix  B.
@@ 3.3 Adaptive reward design
Outputs that match the expected structural format—consisting of <relevance>, <analysis>, and <answer> parts in order—receive a reward of 1; others 0.